OpenStudy (anonymous):

A hardware supplier manufactures three kinds of clamps, types A, B, and C. Production restrictions force it to make 20 more type C clamps than the total of the other types and twice as many type B clamps as type A clamps. The shop must produce 380 clamps per day. How many of each type are made per day?

OpenStudy (anonymous):

A+B+20=C B=2A A+B+C=360 solve the three equtions in three unknowns

OpenStudy (anonymous):

aww A+B+C=380

OpenStudy (anonymous):

Ok I got A=60 B=120 C=200

OpenStudy (anonymous):

A+B+C=380 A+B-C=-20 ------------ 2A+2B=360 B+2B=360 3B=360 B=120

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

Thank you! that helped me so much